# Model Extrapolation Expedites Alignment

### v3

@@ 1 Introduction
Our work is first motivated by the observation that preference alignment training typically does not inject new knowledge into models, thereby likely inducing only small changes of model parameters. We support this hypothesis through three arguments. First, mainstream alignment algorithms like RLHF and DPO incorporate a constraint term (e.g., the KL divergence term) to prevent excessive deviation from the initial SFT checkpoint. Second, in recent open-source LLM alignment projects (Tunstall et al., 2023; Wang et al., 2023; Ivison et al., 2023), preference alignment training usually adopts smaller learning rates (e.g., 5e-7) and fewer training steps (e.g., 400~500 steps) than SFT. Third, we take the zephyr-7b-dpo model (Tunstall et al., 2023) trained by HuggingFace as a specific instance. For any two among the pre-trained, SFT, and DPO checkpoints and for any corresponding parameter tensors $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, we compute the Frobenius norm $\|\mathbf{P}_{1}-\mathbf{P}_{2}\|$ (and a normalized variant)111 The Frobenius norm of tensor $\mathbf{P}$ is defined as: $\|\mathbf{P}\|=\sqrt{\sum_{i}\mathrm{P}_{i}^{2}}$, while the normalized variant is defined as: $\|\mathbf{P}\|=\sqrt{\frac{1}{|\mathbf{P}|}\sum_{i}\mathrm{P}_{i}^{2}}$, where $|\mathbf{P}|$ denotes the element number of $\mathbf{P}$. . In Table 1, we show that the parameter change of alignment training (i.e., from SFT to DPO) is fairly small, whose absolute value of normalized Frobenius distance is merely $6.348\times 10^{-6}$, and is also significantly smaller than that of SFT (i.e., from Pre-trained to SFT). Therefore, in this work we hypothesize that preference alignment training usually involves only small parameter changes.
@@ 2.2 First-order Approximation
Based on the aforementioned observation, we suppose that the parameter change from $\mathcal{M}_{0}$ to $\mathcal{M}_{1}$, denoted as $\|{\bm{\theta}}_{1}-{\bm{\theta}}_{0}\|=\|\Delta{\bm{\theta}}\|$, is small. We can formally perform a Taylor Expansion of $\omega$ at ${\bm{\theta}}_{0}$ and retain the first-order term:
@@
where we define $\gamma\in[0,1]$ to ensure that $\|\gamma\Delta{\bm{\theta}}\|$ remains small. In particular, setting $\gamma=1$ gives:
@@ 2.3 ExPO: Model Extrapolation
In the above first-order approximation, we constrain $\gamma\in[0,1]$ to maintain the approximation’s validity along the straight-line path between ${\bm{\theta}}_{0}$ and ${\bm{\theta}}_{1}$. We now consider extending this approximation to the “extension” of the line connecting ${\bm{\theta}}_{0}$ and ${\bm{\theta}}_{1}$ beyond ${\bm{\theta}}_{1}$. Let $\gamma>1$ and define $\alpha=\gamma-1>0$, denoting ${\bm{\theta}}_{2}={\bm{\theta}}_{0}+\gamma\Delta{\bm{\theta}}={\bm{\theta}}_{0}+(1+\alpha)\Delta{\bm{\theta}}$. By choosing appropriate $\alpha$ such that $\|(1+\alpha)\Delta{\bm{\theta}}\|$ remains small, we can reformulate the first-order approximation as:
@@
From Equation 3, we approximately have $\omega({\bm{\theta}}_{2})>\omega({\bm{\theta}}_{1})$. This suggests that, starting from a partially-aligned model $\mathcal{M}_{1}$ and its SFT checkpoint $\mathcal{M}_{0}$, by selecting appropriate $\alpha>0$, we can construct a new model $\mathcal{M}_{2}$ parameterized by ${\bm{\theta}}_{2}$ through amplifying the parameter change $\Delta{\bm{\theta}}$:
@@ 3.2 Analysis of Varying Training Steps
The optimal $\alpha$ values for $\mathcal{M}_{2}^{10\%}$, $\mathcal{M}_{2}^{20\%}$, $\mathcal{M}_{2}^{40\%}$, and $\mathcal{M}_{2}^{100\%}$ are 8.0, 2.5, 0.5, and 0.3, respectively. Figure 3 illustrates the reward distributions of these extrapolated models, showing that their response distributions shift toward higher reward regions compared to the original $\mathcal{M}_{1}^{*}$ models. In Figure 4, we show that increasing $\alpha$ within a reasonable range consistently improves alignment performance. However, excessively large $\alpha$ causes sharp performance drops and abnormal response length increases (e.g., generating gibberish or failing to terminate). This indicates that overly large $\alpha$ violates the first-order approximation (Equation 4) as $\|(1+\alpha)\Delta{\bm{\theta}}\|$ becomes too large. Additionally, since more training steps lead to larger $\|\Delta{\bm{\theta}}\|$, smaller $\alpha$ values are required for models with more training steps (e.g., $\mathcal{M}_{1}^{100\%}$) to maintain the validity of Equation 4, which is consistent with our hyperparameter search results.
@@ 3.4 Analysis of Training Configurations
We increase the training epochs or learning rate for $\mathcal{M}_{1}$. Table 4 shows that while both adjustments improve $\mathcal{M}_{1}$’s performance, they also reduce the benefits of model extrapolation (lower $\mathcal{M}_{2}$ performance) and yield smaller optimal $\alpha$ values. Meanwhile, the $\mathcal{M}_{1}$ models trained with more epochs or larger learning rates generate significantly longer responses compared to the default setup. This suggest that both adjustments also make models prone to learning the length bias in training data, thereby degrading $\Delta{\bm{\theta}}$’s quality and the gains from ExPO. Notably, when training epochs are set to 3, $\mathcal{M}_{1}$ cannot benefit from ExPO, likely because the first-order approximation (Equation 4) no longer holds as $\|\Delta{\bm{\theta}}\|$ becomes too large.
@@ 4.1 Applying ExPO to More Existing, Already-aligned LLMs
In § 3.2, we observed that ExPO also brings noticeable performance improvements to the fully-trained zephyr-7b-dpo. This motivates us to apply ExPO to more existing, already-aligned LLMs. As hypothesized in § 1, the normally-trained models should also satisfy the first-order approximation premise, i.e., $\|\Delta{\bm{\theta}}\|$ is small. We select twelve open-source models from HuggingFace for experiments (see Appendix C for their model IDs):
@@ 4.3 Discussion on Failure Cases
In both cases, ExPO’s underlying first-order approximation can become invalidated as the resulting $\|\Delta{\bm{\theta}}\|$ is too large. Therefore, we suggest that more deliberate strategies are needed when applying ExPO to models with large parameter changes, e.g., by leveraging the intermediate checkpoints. We note that recent work has made promising exploration (Lin et al., 2025) and expect more follow-up studies in future work.

### v4

@@ 1 Introduction
Our work is first motivated by the observation that preference alignment training typically does not inject new knowledge into models, thereby likely inducing only small changes of model parameters. We support this hypothesis through three arguments. First, mainstream alignment algorithms like RLHF and DPO incorporate a constraint term (e.g., the KL divergence term) to prevent excessive deviation from the initial SFT checkpoint. Second, in recent open-source LLM alignment projects (Tunstall et al., 2023; Wang et al., 2023; Ivison et al., 2023), preference alignment training usually adopts smaller learning rates (e.g., 5e-7) and fewer training steps (e.g., 400~500 steps) than SFT. Third, we take the zephyr-7b-dpo model (Tunstall et al., 2023) trained by HuggingFace as a specific instance. For any two among the pre-trained, SFT, and DPO checkpoints and for any corresponding parameter tensors $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, we compute the Frobenius norm $\left\|\mathbf{P}_{1}-\mathbf{P}_{2}\right\|$ (and a normalized variant)111 The Frobenius norm of tensor $\mathbf{P}$ is defined as: $\left\|\mathbf{P}\right\|=\sqrt{\sum_{i}\mathrm{P}_{i}^{2}}$, while the normalized variant is defined as: $\left\|\mathbf{P}\right\|=\sqrt{\frac{1}{|\mathbf{P}|}\sum_{i}\mathrm{P}_{i}^{2}}$, where $|\mathbf{P}|$ denotes the element number of $\mathbf{P}$. . In Table 1, we show that the parameter change of alignment training (i.e., from SFT to DPO) is fairly small, whose absolute value of normalized Frobenius distance is merely $6.348\times 10^{-6}$, and is also significantly smaller than that of SFT (i.e., from Pre-trained to SFT). Therefore, in this work we hypothesize that preference alignment training usually involves only small parameter changes.
@@ 2.2 First-order Approximation
Based on the aforementioned observation, we suppose that the parameter change from $\mathcal{M}_{0}$ to $\mathcal{M}_{1}$, denoted as $\left\|{\bm{\theta}}_{1}-{\bm{\theta}}_{0}\|=\|\Delta{\bm{\theta}}\right\|$, is small. We can formally perform a Taylor Expansion of $\omega$ at ${\bm{\theta}}_{0}$ and retain the first-order term:
@@
where we define $\gamma\in[0,1]$ to ensure that $\left\|\gamma\Delta{\bm{\theta}}\right\|$ remains small. In particular, setting $\gamma=1$ gives:
@@ 2.3 ExPO: Model Extrapolation
In the above first-order approximation, we constrain $\gamma\in[0,1]$ to maintain the approximation’s validity along the straight-line path between ${\bm{\theta}}_{0}$ and ${\bm{\theta}}_{1}$. We now consider extending this approximation to the “extension” of the line connecting ${\bm{\theta}}_{0}$ and ${\bm{\theta}}_{1}$ beyond ${\bm{\theta}}_{1}$. Let $\gamma>1$ and define $\alpha=\gamma-1>0$, denoting ${\bm{\theta}}_{2}={\bm{\theta}}_{0}+\gamma\Delta{\bm{\theta}}={\bm{\theta}}_{0}+(1+\alpha)\Delta{\bm{\theta}}$. By choosing appropriate $\alpha$ such that $\left\|(1+\alpha)\Delta{\bm{\theta}}\right\|$ remains small, we can reformulate the first-order approximation as:
@@
According to Equation 3, we approximately have $\omega({\bm{\theta}}_{2})>\omega({\bm{\theta}}_{1})$. This suggests that, starting from a partially-aligned model $\mathcal{M}_{1}$ and its SFT checkpoint $\mathcal{M}_{0}$, by selecting appropriate $\alpha>0$, we can construct a new model $\mathcal{M}_{2}$ parameterized by ${\bm{\theta}}_{2}$ through amplifying the parameter change $\Delta{\bm{\theta}}$:
@@ 3.2 Analysis of Varying Training Steps
The optimal $\alpha$ values for $\mathcal{M}_{2}^{10\%}$, $\mathcal{M}_{2}^{20\%}$, $\mathcal{M}_{2}^{40\%}$, and $\mathcal{M}_{2}^{100\%}$ are 8.0, 2.5, 0.5, and 0.3, respectively. Figure 3 illustrates the reward distributions of these extrapolated models, showing that their response distributions shift toward higher reward regions compared to the original $\mathcal{M}_{1}^{*}$ models. In Figure 4, we show that increasing $\alpha$ within a reasonable range consistently improves alignment performance. However, excessively large $\alpha$ causes sharp performance drops and abnormal response length increases (e.g., generating gibberish or failing to terminate). This indicates that overly large $\alpha$ violates the first-order approximation (Equation 4) as $\left\|(1+\alpha)\Delta{\bm{\theta}}\right\|$ becomes too large. Additionally, since more training steps lead to larger $\left\|\Delta{\bm{\theta}}\right\|$, smaller $\alpha$ values are required for models with more training steps (e.g., $\mathcal{M}_{1}^{100\%}$) to maintain the validity of Equation 4, which is consistent with our hyperparameter search results.
@@ 3.4 Analysis of Training Configurations
We increase the training epochs or learning rate for $\mathcal{M}_{1}$. Table 4 shows that while both adjustments improve $\mathcal{M}_{1}$’s performance, they also reduce the benefits of model extrapolation (lower $\mathcal{M}_{2}$ performance) and yield smaller optimal $\alpha$ values. Meanwhile, the $\mathcal{M}_{1}$ models trained with more epochs or larger learning rates generate significantly longer responses compared to the default setup. This suggest that both adjustments also make models prone to learning the length bias in training data, thereby degrading $\Delta{\bm{\theta}}$’s quality and the gains from ExPO. Notably, when training epochs are set to 3, $\mathcal{M}_{1}$ cannot benefit from ExPO, likely because the first-order approximation (Equation 4) no longer holds as $\left\|\Delta{\bm{\theta}}\right\|$ becomes too large.
@@ 4.1 Applying ExPO to More Existing, Already-aligned LLMs
In § 3.2, we observed that ExPO also brings noticeable performance improvements to the fully-trained zephyr-7b-dpo. This motivates us to apply ExPO to more existing, already-aligned LLMs. As hypothesized in § 1, the normally-trained models should also satisfy the first-order approximation premise, i.e., $\left\|\Delta{\bm{\theta}}\right\|$ is small. We select twelve open-source models from HuggingFace for experiments (see Appendix C for their model IDs):
@@ 4.3 Discussion on Failure Cases
In both cases, ExPO’s underlying first-order approximation can become invalidated as the resulting $\left\|\Delta{\bm{\theta}}\right\|$ is too large. Therefore, we suggest that more deliberate strategies are needed when applying ExPO to models with large parameter changes, e.g., by leveraging the intermediate checkpoints. We note that recent work has made promising exploration (Lin et al., 2025) and expect more follow-up studies in future work.
@@ 6 Limitations
Acknowledgements
We thank Sidi Lu, Yufei Tian, Zi-Yi Dou, and other members of the UCLA PlusLab & NLP group as well as anonymous reviewers for their constructive feedback and discussions.
This work was supported by an Amazon AGI Research Award through UCLA-Amazon Science Hub. This work was also supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604) and China Scholarship Council (with No. 202306210211).